Please show work

11. Consider a market with O -240-6P andQ 2P What is the consumer surplus in this market? a) 1,000 b) 300 c) 750 d)500
Answer #1

Answer
the market is in equilibrium at Qd=Qs
240-6P=2P
8P=240
P=30
Q=240-6*30=60
inverse demand curve is
Q=240-6P
6P=240-Q
P=40-(1/6)Q

CS=0.5*(Y-axis coefficient of the demand curve -P)*Q
=0.5*(40-30)*60
=$300

the consumer surplus is $300
option b